# Real-time propagators resummed with nontrivial boundary wavefunctions in a constant electric field

### v1

@@ 7 Conclusion
Interestingly, the action is modified by only the quadratic part and the interaction terms are intact. After all, we have found that the deviation from the standard in-in setup amounts to the self-energy-like terms in the propagators. Although the proper-time representations of the propagators in the modified formalism have been known in the literature, the derivation and the physical meaning in the canonical operator formalism look obscure. Our present derivation based on the functional integration clearly shows that the boundary wavefunction effects are understood as diagrammatic resummation corresponding to the Dyson series. The resulting propagators perfectly coincide with the expressions from the canonical operator formalism, but our careful treatment of the closed-time path in the functional integration formalism has clarified the origin of the additional proper-time contours.
@@ Acknowledgements.
The authors thank Patrick Copinger, Gergely Endrődi, Gergely Markó, and Shi Pu for useful discussions. This work was partially supported by Japan Society for the Promotion of Science (JSPS) KAKENHI Grant No. 22H01216 (K.F.) and No. 24KJ0985 (S.M.).
@@ A.1 Weber parabolic cylinder function
The integral representation of $D_{\mu}(z)$ is known to be
@@ A.2 Explicit expressions in sQED
From the asymptotic forms of the parabolic cylindrical functions, one can read the Bogoliubov coefficients as
@@ A.3 Explicit Expressions in QED
$\sum_{s}\int_{{\bm{p}}}\Bigl[U_{{\bm{p}},s}^{\text{as}}(x^{0},{\bm{x}})\,U_{{\bm{p}},s}^{\text{as}}(x^{0},{\bm{y}})^{\dagger}+V_{{\bm{p}},s}^{\text{as}}(x^{0},{\bm{x}})\,V_{{\bm{p}},s}^{\text{as}}(x^{0},{\bm{y}})^{\dagger}\Bigr]=\delta_{ss^{\prime}}\,\delta^{(3)}({\bm{x}}-{\bm{y}})$ (113)
@@
To construct a complete set for the Dirac equation, one can find the energy dispersion relation from the squared Dirac operator in momentum space, i.e.,
@@
Appendix B Derivation of $D^{--}(x,y)$ and $S^{--}(x,y)$
@@
so we expand the right-hand side as a Dyson series and evaluate each term explicitly.
@@
Then, Eq. (138) yields
@@ Appendix C Derivation of the proper-time representation
C.1 Integral forms of the Weber function
@@ C.2 Propagators in sQED
For latter convenience, we rewrite these equations by using the following identities that hold when the integrals are treated as the principal-value integral, i.e.,
@@
The integration contours are shown in Fig. 4. It is easy to prove Eqs. (168) and (169) by first shrinking the integration contours to paths near the singular points as explicitly indicated in Fig. 4. Then, we see that the contributions from the semicircles are exponentially suppressed depending on the sign of $z_{3}^{2}-z_{0}^{2}$. Using Eqs. (168) and (169), we readily obtain
@@ C.3 Propagators in QED
where the definition of $g(s;x,y)$ is given above Eq. (3). The relative sign difference between $S^{>}(x,y)$ and $S^{<}(x,y)$ is a consequence of the anticommutative nature of fermions. By using the following equations,
@@
$\theta(z_{0}^{2}-z_{3}^{2})\int\limits_{\Gamma_{\mathrm{a}}-\Gamma_{2}-\Gamma_{3}}\mathrm{d}s\,g(s;x,y)=0\,$, (177)
which are analogous to Eqs. (168) and (169). Then, we can rewrite the propagators as

### v3

@@ 7 Conclusion
Interestingly, the action is modified by only the quadratic part and the interaction terms are intact. After all, we have found that the deviation from the standard in-in setup amounts to the self-energy-like terms in the propagators. Although the proper-time representations of the propagators in the modified formalism have been known in the literature, the derivation and the physical meaning in the canonical operator formalism look obscure. Our present derivation based on the functional integration clearly shows that the boundary-wavefunction effects are understood as diagrammatic resummation corresponding to the Dyson series. The resulting propagators perfectly coincide with the expressions from the canonical operator formalism, but our careful treatment of the closed-time path in the functional integration formalism has clarified the origin of the additional proper-time contours.
@@ Acknowledgements.
The authors would like to give special thanks to Patrick Copinger and Shi Pu. In the early stage of this work, the authors communicated with them, and the present work was strongly inspired through discussions with them. Their approach is based on the worldline formalism (see Ref. Copinger:2025ovz ), while our method does not rely on the worldline but employs the closed-time-path formalism. The authors also thank Gergely Endrődi and Gergely Markó for useful discussions. This work was partially supported by Japan Society for the Promotion of Science (JSPS) KAKENHI Grant No. 22H01216 (K.F.) and No. 24KJ0985 (S.M.).
@@ A.1 Weber parabolic cylinder function
The integral representation of $D_{\nu}(z)$ is known to be
@@ A.2 Explicit expressions in sQED
From the asymptotic forms of the Weber parabolic cylinder functions, one can read the Bogoliubov coefficients as
@@ A.3 Explicit expressions in QED
$\sum_{s}\int_{{\bm{p}}}\Bigl[U_{{\bm{p}},s}^{\text{as}}(x^{0},{\bm{x}})\,U_{{\bm{p}},s}^{\text{as}}(x^{0},{\bm{y}})^{\dagger}+V_{{\bm{p}},s}^{\text{as}}(x^{0},{\bm{x}})\,V_{{\bm{p}},s}^{\text{as}}(x^{0},{\bm{y}})^{\dagger}\Bigr]=\,\delta^{(3)}({\bm{x}}-{\bm{y}})$ (113)
@@
To obtain the explicit solutions, it is useful to consider the squared Dirac operator:
@@
Appendix B Explicit calculations of the Dyson series for $D^{--}(x,y)$ and $S^{--}(x,y)$
@@
so we expand the right-hand side as the Dyson series and evaluate each term explicitly.
@@
In the last equality, we use Eq.(105). Then, Eq. (138) yields
@@ Appendix C Derivation of the proper-time representation
C.1 $\mathscr{I}^{(\mp)}_{\nu}$ and $\mathscr{I}^{\rm a}_{\nu}$
@@ C.2 Propagators in sQED
For latter convenience, we rewrite these equations by using the following identities that hold when the integrals are treated as the principal-value integral and when the integration contours are shrunk to paths near the singular points as explicitly indicated in Fig. 4, i.e.,
@@
The integration contours are shown in Fig. 4. We can readily see that the contributions from the semicircles are exponentially suppressed depending on the sign of $z_{3}^{2}-z_{0}^{2}$, and the contribution from a part of the contour parallel to the real axis also vanishes in the infinitesimal limit of the contour. Using Eqs. (168) and (169), we obtain
@@ C.3 Propagators in QED
where the definition of $g(s;x,y)$ is given above Eq. (3). The relative sign difference between $S^{>}(x,y)$ and $S^{<}(x,y)$ is a consequence of the anticommutative nature of fermions. By using the following equations that are valid under the same conditions for Eqs. (168) and (169),
@@
$\theta(z_{0}^{2}-z_{3}^{2})\int\limits_{\Gamma_{\mathrm{a}}-\Gamma_{2}-\Gamma_{3}}\mathrm{d}s\,g(s;x,y)=0\,.$ (177)
The integration contours are shown in Fig. 4. Then, we can rewrite the propagators as